OpenStudy (anonymous):

Using implicit differentiation find dy/dx: 1+x=sin(xy²)

OpenStudy (anonymous):

do you have anything so far?

OpenStudy (anonymous):

1= cos(xy²)(xyy'+y²) ?

OpenStudy (anonymous):

\[\huge 1+x=sin(xy^2) \] \[\huge [1+x]'=[sin(xy^2)]' \] \[\huge 1=cos(xy^2)\cdot (xy^2)' \] can you do that derivative on the right?

OpenStudy (anonymous):

almost..

OpenStudy (anonymous):

x2yy'+y² ?

OpenStudy (anonymous):

\[\huge (xy^2)'=(x)'y^2+x(y^2)' \] \[\huge (xy^2)'=1\cdot y^2+x\cdot 2yy' \]

OpenStudy (anonymous):

yep.. you got it...:)

OpenStudy (anonymous):

so we have... \[\huge 1=cos(xy^2)\cdot (y^2+2xyy') \]

OpenStudy (anonymous):

solve for y'...

OpenStudy (anonymous):

:/ I don't know how. I get: 1/cos(xy²)= (y2+2xyy′) then Idk

OpenStudy (anonymous):